# Kino-PAX: Highly Parallel Kinodynamic Sampling-based Planner

### v1

@@ I Introduction
Autonomous robotic systems are increasingly deployed in dynamic environments, requiring fast, reactive motion planning that accounts for the robot’s complex kinematics and dynamics. Solving the kinodynamic motion planning problem quickly is critical for ensuring both functionality and safety. Sampling-based motion planners (SBMPs) have proven effective for various difficult problems, such as complex dynamics [1, 2, 3, 4, 5], complex tasks [6, 7, 8], and stochastic dynamics [9, 10, 11, 12, 13, 14]. Nevertheless, they are typically designed for serial computation, limiting their speed to CPU clock rate. While recent methods can find solutions within seconds for simple systems and tens of seconds for complex ones [1, 2, 3], this is insufficient for real-time reactivity. Given the plateau in improvements to serial computation and CPU clock speeds, parallel devices like GPUs offer promising speedups. However, current SBMP algorithms are inherently sequential and inefficient when parallelized. In this work, we aim to enable real-time motion planning for complex and high-dimensional kinodynamical systems by exploiting the parallel architecture of GPU-like devices.
@@
In summary, our contributions are four-fold: (i) Kino-PAX, a highly parallel kinodynamic SBMP designed to leverage the parallel architecture of GPU-like devices, (ii) discussion on efficient GPU implementation of Kino-PAX, (iii) a thorough analysis and proof of probabilistic completeness, and (iv) benchmarks showing the efficiency and efficacy of Kino-PAX for complex and high-dimensional dynamical systems. Our results show that Kino-PAX achieves up to three-orders-of-magnitude improvement in computation time compared to our baselines, which use CPU parallelization. In all evaluated problems, Kino-PAX finds solutions in the order of $10$ milliseconds, representing significant progress in enabling real-time kinodynamic motion planning.
@@ I-A Related Work
SBMPs have a long-standing history in addressing the geometric motion planning problem, as established by foundational works such as Probabilistic RoadMaps (PRM) [15], Expansive-Space Tree (EST) [16], Rapidly-exploring Random Tree (RRT) [17], etc. In general terms, SBMP techniques involve finding a path from a starting configuration to a goal region by constructing a graph or tree, where nodes represent geometric configuration and straight line edges represent transitions in the configuration space. Traditionally, these algorithms operate serially on CPU devices. However, the increasing demand for rapid replanning for complex systems in unknown and dynamic environments has driven the development of parallelization methods for geometric SBMPs. These parallelized approaches have been applied to both CPU-based planners [18, 19] and GPU-based implementations [20].
Work [18] decomposes fundamental operations of geometric SBMPs, such as collision checking, forward kinematics, and nearest neighbor search, into unconventional data layouts. This approach enables parallelism without requiring specialized hardware. Work [19] proposes growing multiple trees in parallel, where information is shared between trees, improving visibility of the search space. In [20], the authors leverage GPUs to solve the geometric path planning problem by adapting FMT* [21] for parallelized graph search, achieving orders of magnitude performance improvements over its serial counterpart. While these algorithms have demonstrated orders of magnitude improvement over serial implementations, they are strictly limited to geometric problems.
@@
In contrast to the parallelized geometric planning solutions discussed above, parallelization for planning under the constraints of general dynamical systems is relatively understudied in the field. An approach to parallelization for the kinodynamic problem is a coarse-grained method, where multiple trees of classical sampling-based motion planners are generated in parallel, and the first solution found is returned [22]. This technique improves average-case performance [23], and is employed in Section VI to perform CPU-based parallelization as baseline. However, the inherent sequential nature of these motion planners make them inefficient for massive parallelization. In this work, we propose a novel algorithm that enables efficient application to parallel devices.
@@ II Problem Formulation
Consider a robotic system operating within a bounded workspace $W\subset\mathbb{R}^{d}$, where $d\in\{2,3\}$. This workspace contains a finite set of obstacles $\mathcal{O}$, where each obstacle $o\in\mathcal{O}$ is a closed subset of $W$, i.e., $o\subseteq W$. The dynamics of the robot’s motion is given by
@@
In addition to motion constraints defined by the dynamics in (1) and obstacles in $\mathcal{O}$, we consider state constraints, e.g., bound on the velocity. To this end, we define the set of valid states, i.e., states at which the robot does not violate its state constraints and does not collide with an obstacle, as the constraint-free set and denote it by $X_{\text{free}}\subseteq X$. Then, given initial state $x_{\text{init}}\in X_{\text{free}}$, time duration $t_{f}\geq 0$, and control trajectory $\mathbf{u}:[0,t_{f}]\to U$, a state trajectory $\mathbf{x}:[0,t_{f}]\to X$ is induced, where
@@
Trajectory $\mathbf{x}$ is called valid if, $\forall t\in[0,t_{f}]$, $x(t)\in X_{\text{free}}$.
In motion planning, the interest is to find a valid trajectory $\mathbf{x}$ that visits a given goal set $X_{\text{goal}}\subseteq X_{\text{free}}$. Therefore, by following this trajectory, the robot is able to respect all of its motion (kinodynamic) constraints, avoid collisions with obstacles, and reach its goal. In this work, we focus on kinodynamic motion planning with an emphasis on computational efficiency through parallelism.
@@
Consider a robot with dynamics in (1) in workspace $W$ consisting of obstacle set $\mathcal{O}$. Given an initial state $x_{\text{init}}\in X_{\text{free}}\subseteq X$ and goal region $X_{\text{goal}}\subseteq X_{\text{free}}$, efficiently find a control trajectory $\mathbf{u}:[0,t_{f}]\to U$ such that its induced trajectory $\mathbf{x}$ through (2) is valid and reaches goal, i.e., $\mathbf{x}(0)=x_{\text{init}}$ and
$\mathbf{x}(t)$ $\in X_{\text{free}}$ $\forall t\in[0,t_{f}],$
@@ III Kino-PAX
Figure 1: Overview of Kino-PAX.
(a) $V_{E}$ for extension
(b) Extend $V_{E}$ nodes
(c) Update $V_{U}$
(d) Update $V_{E}$ and $V_{O}$
Figure 2: Illustration of a single iteration in the Kino-PAX expansion process. (a) The current sets $V_{E}$ and $V_{O}$, where the numbers in each grid cell represent the value of $P_{accept}(\mathcal{R}_{i})$. (b) Expansion of each node in $V_{E}$ with a branching factor of $\lambda=2$. (c) Acceptance of promising samples and the update of $P_{accept}(\mathcal{R}_{i})$. (d) The updated set $V_{E}$, prepared for expansion in the next iteration.
@@
With these principles in mind, we introduce Kinodynamic Parallel Accelerated eXpansion (Kino-PAX), a highly parallel kinodynamic SBMP. An overview of Kino-PAX is shown in Fig. 1. Kino-PAX grows a tree of trajectory segments in parallel. This is achieved by decomposing the iterative tree growth process, i.e., selection of nodes, extension, validity checking, and adding new nodes to the tree, into three massively parallel subroutines. Each subroutine follows the key principles of thread independence, balanced workloads, and low-latency memory utilization. Additionally, to ensure fast and efficient planning iterations, we minimize the communication needed in the synchronization steps between subroutines, e.g., CPU-GPU communication.
@@ 13                   Increment $n_{invalid}(\mathcal{R}_{i})$ ;
After initialization, the main loop of the algorithm begins (Lines 4–8). In each iteration, the Propagate (Alg. 2) function is called to propagate the set $V_{E}$ in parallel (Line 5, Fig. 2(b)). Each node $x\in V_{E}$ is expanded $\lambda\in\mathbb{N^{+}}$ times using $\lambda$ threads, where each thread handles one expansion of $x$ (Alg. 2, Lines 1–2). For each thread, a control $u\in U$ and a time duration $dt\in(0,T_{prop}]$ are randomly sampled, and the node’s continuous state $x$ is propagated using dynamics in (1) to generate a new sample state $x^{\prime}$ (Alg. 2, Lines 3–4). Next, the corresponding region of $x^{\prime}$, $\mathcal{R}^{k}_{i}$, is calculated (Alg. 2, Line 5). Then, in Line 6, the trajectory segment from $x$ to $x^{\prime}$ is checked for validity, i.e., if it is in $X_{\text{free}}$. Throughout the trajectory segment, a user-defined collision check is performed (our implementation uses a coarse-phase bounding volume hierarchies method discussed in [20]). If the extended segment is valid, we increment the total number of valid samples in $\mathcal{R}_{i}$. Next, $x^{\prime}$ is added to the set $V_{U}$ if its corresponding region $\mathcal{R}^{k}_{i}$ is unvisited; if $\mathcal{R}^{k}_{i}$ already contains a node, then $x^{\prime}$ is added to $V_{U}$ with probability $P_{accept}(\mathcal{R}_{i})$, which favors promising samples (Alg. 2, Lines 8-9, Fig. 2(c)). Alternatively, if the trajectory segment is invalid, we increment the count of invalid samples in $\mathcal{R}_{i}$, as shown in Line 11. This information guides future propagation iterations away from regions that are frequently invalid, improving search efficiency.
@@ III-A3 Node Selection
After all samples in $V_{E}$ have been expanded, the UpdateEstimates (Alg. 3) subroutine is called. In this subroutine, metrics for each visited region (i.e., a region with a node $x\in\mathcal{T}$), denoted by $\mathcal{R}_{avail}$, are updated in parallel, with a thread handling a unique region $\mathcal{R}_{i}\in\mathcal{R}_{avail}$, calculating $Cov(\mathcal{R}_{i})$ and $FreeVol(\mathcal{R}_{i})$ (Alg. 3, Lines 1-3). For each thread, $Cov(\mathcal{R}_{i})$ is set to the number of visited sub-regions within $\mathcal{R}_{i}$, and $FreeVol(\mathcal{R}_{i})$ is calculated as
@@
Once all score values have been updated, each visited region’s $P_{accept}(\mathcal{R}_{i})$ probability is refined (Alg. 3 Lines 5-6, Fig. 2(c)). This process, again, is done in parallel with a thread being designated to a unique $\mathcal{R}_{i}\in\mathcal{R}_{avail}$ and $P_{accept}(\mathcal{R}_{i})$ being set by
@@
Once the $P_{accept}(\mathcal{R}_{i})$ probabilities have been updated for all available regions, the UpdateNodeSets subroutine is called (Alg. 1, Line $9$, Fig. 2(d)). In Lines $1-3$ of Alg. 4, we remove samples from the expansion set $V_{E}$ randomly with probability $1-P_{accept}(\mathcal{R}_{i})$, ensuring that promising samples remain in $V_{E}$. Then, we move the newly generated samples, $V_{U}$, to $\mathcal{T}$ and add them to the expansion set $V_{E}$. If any newly generated nodes satisfy goal criteria, we return the valid trajectory $\mathbf{x}$ (Alg. 4, Lines 4-6). Finally, we move inactive samples in $V_{O}$ to the expansion set if deemed promising with the updated search information (Alg. 4, Lines 7-9).
@@ IV-A Tuning Parameter
where $\lambda_{max}$ is the user-defined maximum branching factor and $|\mathcal{T}|$ and $|V_{E}|$ are the numbers of nodes in $\mathcal{T}$ and $V_{E}$, respectively. Eq. (6) ensures that in the early iterations, when $|\mathcal{T}|$ is much smaller than $t_{e}$, a larger $\lambda$ is used. This better uses available throughput and helps accelerate the initial stages of tree propagation. As $|\mathcal{T}|$ approaches $t_{e}$ and $|V_{E}|$ is large, a smaller $\lambda$ is used to prevent device serialization and increase tree sparsity.
@@
IV-B Efficient Application to Parallel Devices
Kino-PAX’s propagation subroutine implementation is well-suited for parallelism due to three main factors. First, during node expansion, we utilize low-latency on-chip memory to distribute the state $x\in V_{E}$ among all $\lambda$ threads assigned to expand the node. Second, we minimize thread divergence by having each thread create only a single trajectory segment. Third, the acceptance of new samples and their addition to $V_{U}$ is performed completely independently, using (5) and unique thread identifiers to avoid memory write conflicts.

### v2

@@ Abstract
Index Terms:
Constrained Motion Planning, Motion and Path Planning, Computer Architecture for Robotic and Automation.
@@ I Introduction
Autonomous robotic systems are increasingly deployed in dynamic environments, requiring fast, reactive motion planning that accounts for the robot’s complex kinematics and dynamics. Solving the kinodynamic motion planning problem quickly is critical for ensuring both functionality and safety. Sampling-based motion planners (SBMPs) have proven effective for various difficult problems, such as complex dynamics [1, 2, 3, 4, 5], complex tasks [6, 7, 8], and stochastic dynamics [9, 10, 11]. Nevertheless, they are typically designed for serial computation, limiting their speed to CPU clock rate. While recent methods can find solutions within seconds for simple systems and tens of seconds for complex ones [1, 2, 3], this is insufficient for real-time reactivity. Given the plateau in improvements to serial computation and CPU clock speeds, parallel devices like GPUs offer promising speedups. However, current SBMP algorithms are inherently sequential and inefficient when parallelized. In this work, we aim to enable real-time motion planning for complex and high-dimensional kinodynamical systems by exploiting the parallel architecture of GPU-like devices.
@@
In summary, our contributions are four-fold: (i) Kino-PAX, a highly parallel kinodynamic SBMP designed to leverage the parallel architecture of GPU-like devices, (ii) a discussion of Kino-PAX’s hyperparameters and its efficient application to highly parallel devices, (iii) a thorough analysis and proof of probabilistic completeness, and (iv) benchmarks showing the efficiency and efficacy of Kino-PAX for complex and high-dimensional dynamical systems. Our results show that Kino-PAX achieves up to three-orders-of-magnitude improvement in computation time compared to our baselines, which use CPU parallelization. In all evaluated problems, Kino-PAX finds solutions in the order of $10$ milliseconds, representing significant progress in enabling real-time kinodynamic motion planning.
@@ I-A Related Work
SBMPs have a long-standing history in addressing the geometric motion planning problem, as established by foundational works such as Probabilistic RoadMaps (PRM) [12], Expansive-Space Tree (EST) [13], Rapidly-exploring Random Tree (RRT) [14], etc. In general terms, SBMP techniques involve finding a path from a starting configuration to a goal region by constructing a graph or tree, where nodes represent geometric configurations and straight line edges represent transitions in the configuration space. Traditionally, these algorithms operate serially on CPU devices. However, the increasing demand for rapid replanning for complex systems in unknown and dynamic environments has driven the development of parallelization methods for geometric SBMPs. These parallelized approaches have been applied to both CPU-based planners [15, 16, 17, 18, 19, 20] and GPU-based implementations [21].
The CPU-based methods in [15, 16, 17] use coarse-grained parallelization, where threads independently construct trees and exchange search information. These methods offer straightforward implementation by leveraging existing serial SBMPs but they achieve limited improvements in planning rates, and are not applicable to many-core devices. More similar to our work, the approaches in [18, 19, 20] focus on fine-grained parallelization to accelerate the construction of a single tree. Works [18, 19] introduce parallel RRT and RRT* methods that leverage thread-safe atomic operations and a novel concurrent data structure for efficient nearest neighbor search. In contrast, [20] decomposes core operations of geometric SBMPs into unconventional data layouts that enable parallelism without specialized hardware. While these works present promising fine-grained parallelization techniques, they are not easily adaptable to kinodynamic planning and are unsuitable for many-core technology due to high inter-thread communication costs.
Most similar to our work, [21] leverages GPUs to solve the geometric path planning problem by adapting FMT* [22] for parallelized graph search. Their implementation achieves orders-of-magnitude performance improvements over its serial counterpart; however, it is strictly limited to geometric problems.
@@
In contrast to the parallelized geometric planning solutions discussed above, parallelization for planning under the constraints of general dynamical systems is relatively understudied in the field. An approach to parallelization for the kinodynamic problem is a coarse-grained method, where multiple trees of classical SBMPs are generated in parallel, and the first solution found is returned [23]. This technique improves average-case performance [24], and is employed in Section VI to perform CPU-based parallelization as a baseline. However, the inherent sequential nature of these motion planners make them inefficient for massive parallelization. In this work, we propose a novel algorithm that enables efficient application to highly parallel devices.
@@ II Problem Formulation
Consider a robotic system operating within a bounded workspace $W\subset\mathbb{R}^{d}$, where $d\in\{2,3\}$. This workspace contains a finite set of obstacles $\mathcal{O}$, where each obstacle $o\in\mathcal{O}$ is a closed subset of $W$, i.e., $o\subset W$. The dynamics of the robot’s motion is given by
@@
In addition to motion constraints defined by the dynamics in (1) and obstacles in $\mathcal{O}$, we consider state constraints, e.g., bound on the velocity. To this end, we define the set of valid states, i.e., states at which the robot does not violate its state constraints and does not collide with an obstacle, as the valid set and denote it by ${X_{\text{valid}}}\subseteq X$. Then, given initial state $x_{\text{init}}\in{X_{\text{valid}}}$, time duration $t_{f}\geq 0$, and control trajectory $\mathbf{u}:[0,t_{f}]\to U$, a state trajectory $\mathbf{x}:[0,t_{f}]\to X$ is induced, where
@@
Trajectory $\mathbf{x}$ is called valid if, $\forall t\in[0,t_{f}]$, $x(t)\in{X_{\text{valid}}}$.
In motion planning, the interest is to find a valid trajectory $\mathbf{x}$ that visits a given goal set $X_{\text{goal}}\subseteq{X_{\text{valid}}}$. Therefore, by following this trajectory, the robot is able to respect all of its motion (kinodynamic) constraints, avoid collisions with obstacles, and reach its goal. In this work, we focus on kinodynamic motion planning with an emphasis on computational efficiency through parallelism.
@@
Consider a robot with dynamics in (1) in workspace $W$ consisting of obstacle set $\mathcal{O}$. Given an initial state $x_{\text{init}}\in{X_{\text{valid}}}\subseteq X$ and goal region $X_{\text{goal}}\subseteq{X_{\text{valid}}}$, efficiently find a control trajectory $\mathbf{u}:[0,t_{f}]\to U$ such that its induced trajectory $\mathbf{x}$ through (2) is valid and reaches goal, i.e., $\mathbf{x}(0)=x_{\text{init}}$ and
$\mathbf{x}(t)$ $\in{X_{\text{valid}}}$ $\forall t\in[0,t_{f}],$
@@ III Kino-PAX
Figure 1: Illustration of the Kino-PAX expansion process: (a) The current sets $V_{E}$ and $V_{O}$, where the numbers in each grid cell represent the value of $P_{accept}(\mathcal{R}_{i})$. (b) Expansion of $V_{E}$ with branching factor $\lambda=2$. (c) Acceptance of promising samples and update of $P_{\text{accept}}(\mathcal{R}_{i})$. (d) Updated $V_{E}$, ready for the next iteration. (e) Expansion of $V_{E}$, producing a valid trajectory to $X_{\text{goal}}$.
@@
With these principles in mind, we introduce Kinodynamic Parallel Accelerated eXpansion (Kino-PAX), a highly parallel kinodynamic SBMP. Kino-PAX grows a tree of trajectory segments in parallel. This is achieved by decomposing the iterative tree growth process, i.e., selection of nodes, extension, validity checking, and adding new nodes to the tree, into three massively parallel subroutines. Each subroutine follows the key principles of thread independence, balanced workloads, and low-latency memory utilization. Additionally, to ensure fast and efficient planning iterations, we minimize the communication needed in the synchronization steps between subroutines, e.g., CPU-GPU communication.
@@ 13                   Increment $n_{invalid}(\mathcal{R}_{i})$ ;
After initialization, the main loop of the algorithm begins (Alg. 1, Lines 4–8). In each iteration, the Propagate (Alg. 2) function is called to propagate the set $V_{E}$ in parallel (Alg. 1, Line 5, Fig. 1(b)). Each node $x\in V_{E}$ is expanded $\lambda\in\mathbb{N^{+}}$ times using $\lambda$ threads, where each thread handles one expansion of $x$ (Alg. 2, Lines 1–2). For each thread, a control $u\in U$ and a time duration $dt\in(0,T_{prop}]$, where $T_{prop}$ is a user-defined constant that sets the maximum propagation time, are randomly sampled, and the node’s continuous state $x$ is propagated using dynamics in (1) to generate a new sample state $x^{\prime}$ (Alg. 2, Lines 3–4). Next, the corresponding region of $x^{\prime}$, $\mathcal{R}^{k}_{i}$, is calculated (Alg. 2, Line 5). Then, in Line 6, the trajectory segment from $x$ to $x^{\prime}$ is checked for validity, i.e., if it is in $X_{\text{valid}}$. Throughout the trajectory segment, a user-defined collision check is performed (our implementation uses a coarse-phase bounding volume hierarchies method discussed in [21]). If the extended segment is valid, we increment the total number of valid samples in $\mathcal{R}_{i}$ (Alg. 2, Line 7). Next, $x^{\prime}$ is added to the set $V_{U}$ if its corresponding region $\mathcal{R}^{k}_{i}$ is unvisited; if $\mathcal{R}^{k}_{i}$ already contains a node, then $x^{\prime}$ is added to $V_{U}$ with probability $P_{accept}(\mathcal{R}_{i})$, which favors promising samples (Alg. 2, Lines 8-9, Fig. 1(c)). Alternatively, if the trajectory segment is invalid, we increment the count of invalid samples in $\mathcal{R}_{i}$, as shown in Line 11. This information guides future propagation iterations away from regions that are frequently invalid, improving search efficiency.
@@ III-A3 Node Selection
After all samples in $V_{E}$ have been expanded, the UpdateEstimates (Alg. 3) subroutine is called (Alg. 1, Line 6). In this subroutine, metrics for each visited region (i.e., a region with a node $x\in\mathcal{T}$), denoted by $\mathcal{R}_{avail}$, are updated in parallel, with a thread handling a unique region $\mathcal{R}_{i}\in\mathcal{R}_{avail}$, calculating $Cov(\mathcal{R}_{i})$ and $FreeVol(\mathcal{R}_{i})$ (Alg. 3, Lines 1-3). For each thread, $Cov(\mathcal{R}_{i})$ is set to the number of visited sub-regions within $\mathcal{R}_{i}$, and $FreeVol(\mathcal{R}_{i})$ is calculated as
@@
Once all score values have been updated, each visited region’s $P_{accept}(\mathcal{R}_{i})$ probability is refined (Alg. 3, Lines 5-6, Fig. 1(c)). This process, again, is done in parallel with a thread being designated to a unique $\mathcal{R}_{i}\in\mathcal{R}_{avail}$ and $P_{accept}(\mathcal{R}_{i})$ being set by
@@
Once the $P_{accept}(\mathcal{R}_{i})$ probabilities have been updated for all available regions, the UpdateNodeSets subroutine is called (Alg. 1, Line $7$, Fig. 1(d)). In Lines $1-3$ of Alg. 4, we remove samples from the expansion set $V_{E}$ randomly with probability $1-P_{accept}(\mathcal{R}_{i})$, ensuring that promising samples remain in $V_{E}$. Then, we move the newly generated samples, $V_{U}$, to $\mathcal{T}$ and add them to the expansion set $V_{E}$. If any newly generated nodes satisfy goal criteria, we return the valid trajectory $\mathbf{x}$ (Alg. 4, Lines 4-6). Finally, we move inactive samples in $V_{O}$ to the expansion set if deemed promising with the updated search information (Alg. 4, Lines 7-9).
@@ IV-A Tuning Parameter
where $\lambda_{max}$ is the user-defined maximum branching factor and $|\mathcal{T}|$ and $|V_{E}|$ are the numbers of nodes in $\mathcal{T}$ and $V_{E}$, respectively. Eq. (6) ensures that in the early iterations, when $|\mathcal{T}|$ is much smaller than $t_{e}$, a larger $\lambda$ is used. This approach effectively uses available throughput and accelerates the initial stages of tree propagation. As $|\mathcal{T}|$ approaches $t_{e}$ and $|V_{E}|$ grows large, a smaller $\lambda$ is used to stabilize the growth rate of $\mathcal{T}$.
@@
IV-B Decomposition Tuning
The search efficiency of Kino-PAX is dependent on the choice of space decomposition. An improper decomposition, e.g., one that is too coarse or too fine, can lead to inefficiencies. A decomposition that is too fine may cause slower updates due to the large number of regions. On the other hand, a decomposition that is too coarse may lead to poor approximation of promising space regions. For instance, if a region frequently generates invalid trajectories but contains critical space for finding a solution, Kino-PAX may experience inefficient search. A method to mitigate this is to use a free-space-obeying decomposition, as proposed in [3]. Nonetheless, Kino-PAX remains probabilistically complete with any valid decomposition.
IV-C Efficient Application to Highly Parallel Devices
Kino-PAX’s propagation subroutine implementation is well-suited for parallelism due to three main factors. First, we utilize low-latency memory by distributing commonly used data to groups of nearby threads. Specifically, the state $x\in V_{E}$ is shared via on-chip memory to all $\lambda$ threads assigned to expand the node. Second, we balance workloads across threads by having each thread create a single trajectory segment, minimizing thread divergence—i.e., variations in execution paths that force threads into serial execution. Third, the acceptance of new samples and their addition to $V_{U}$ is achieved with thread independence, using (5) and unique thread identifiers.